OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4).

OpenStudy (anonymous):

Hint: Point-Slope form: y - y1 = m(x - x1) m = slope Then convert to slope-intercept form: y = mx + b

OpenStudy (anonymous):

Does that help? Or would you like me to go over the steps?

OpenStudy (anonymous):

Go over steps please!

OpenStudy (anonymous):

Okay (: You have a point and the slope so you can use the Point-Slope form. Fill in the x1 and y1 with the corresponding values and remember m = slope: y - y1 = m(x - x1) y - (-4) = -1/3(x - 6) y + 4 = -1/3(x - 6) Do you undertsand so far?

OpenStudy (anonymous):

Yeah somewhat

OpenStudy (anonymous):

do i solve y + 4 = -1/3(x - 6)?

OpenStudy (anonymous):

Yes, first distribute the -1/3. By the way, the x1 and y1 simply mark where you put the x- and y-values from the point. (x, y) (6, -4) Also, when you have two - signs you change them into +

OpenStudy (anonymous):

y = -1/3x +2 ?

OpenStudy (anonymous):

What happened to the + 4 on the left side of the equation?

OpenStudy (anonymous):

You have everyting else right though... y + 4 = -1/3x + 2 Now subtract 4 from both sides of the equation.

OpenStudy (anonymous):

We want to change this into y=mx+b form so we need to get the y alone.